1. Scuffy has the following product info:

sales price 7.50
variable cost 2.25
fixed cost 10,000
units sold 20,000

what is break even in point in sales?

2.Huts sells hot dog $2/each. The variable cost is $1 and $35 is fixed overhead cost. A summer camp wishes to buy 100 hot dogs for $1.25/each. What is the profit for hut?

3.

. Which of the following organization would be most likely to adopt a process costing system?     …...

a. customer homebuilder

b. law office

c. paper manufacture

d. dental office

e. TV sale and services organization

4.

What type of their special short run decision is most likely to be needed to make:

a.       Accept reject a special order

b.      Bring standard product

c.       Make it yourself or buy it from out side

d.      Sell now or process future

e.       All

5.

  1. Short run decisions is management likely to be
  1. accept of reject speed order
  2. pricing standard products
  3. make it yourself buy from outside
  4. sell now or proceed further
  5. all

6.

The discount rate for use in capital budgeting decision is also referred to as

a. a cost of capital

b. the cost of capital

c. the hurdle rate

d. the minimum required rate of return

e. all none

7.

Sales price                     $7.50 per unit

Variable cost                  $2.25 per unit

Fixed cost                      $10,000

Units sold                      20,000

what is net income?
Answer #1

SOLUTION

Question 1

Contribution margin per unit = Sales price - Variable cost

= $7.50 - $2.25 = $5.25

Contribution margin ratio = Contribution margin per unit / Sales price

= $5.25 / 7.50 = 70%

Break even point = Fixed costs / Contribution margin per unit

= $10,000 / 5.25 = 1,905

Break even point in sales = Fixed costs / Contribution margin ratio

= 10,000 / 70% = $14,286

** All the questions provided are independent to each other, so as per HOMEWORKLIB RULES, I have answered first question.